# Diff-PIC: Revolutionizing Particle-In-Cell Simulation for Advancing Nuclear Fusion with Diffusion Models

### v2

@@ 3.2 Physically-Informed Condition Encoder
PIC simulations employ diverse physical conditions as inputs, which necessitate a seamless and continuous transition in the resulting synthesized data as these input conditions vary. Therefore, the condition encoder plays a crucial role, responsible for transforming domain-specific inputs into embeddings comprehensible by the model. These inputs comprise the simulation conditions $\theta$ and the target simulation time step $t_{as}$ that the conditional diffusion model aims to generate.
@@
Given the extensive potential range of simulation conditions and the limited data available during training, an optimal encoder must excel in both interpolation and extrapolation — critical measures of the model’s generalizability. Recall that interpolation capability refers to the encoder’s proficiency in generating suitable embeddings for new conditions that, although not encountered during training, lie between two observed conditions. Extrapolation capability, conversely, pertains to generating embeddings for conditions that fall outside the range of those observed during training. Both capabilities are indispensable for addressing the LPI problem.
@@
To meet the two essentials, we introduce a Physically-Informed Condition Encoder (PICE) to encode the simulation conditions $\theta$, incorporating two distinct types of encoders. For interpolation, we employ Positional Encoding [65] to learn a continuous representation of input conditions, enabling smooth transitions between observed conditions. To augment extrapolation capability, we enhance the encoder with a polynomial architecture, utilizing transformation functions constructed as a linear combination of polynomial basis functions $f_{i}(\theta)$ of varying degrees:
@@
where $n$ denotes the maximum degree, and $\mathcal{P}(\theta)$ can be chosen as Chebyshev polynomials, or Legendre polynomials, based on the characteristics of the condition space. This polynomial enhancement allows the encoder to generate plausible embeddings for conditions well beyond those encountered during training, ensuring robust performance across a broader spectrum of simulation scenarios. Subsequently, we concatenate the embeddings from these two encoders and apply a Multi-Layer Perceptron (MLP) to further refine the embeddings. The MLP, with its trainable parameters, learns to combine and transform the concatenated embeddings, resulting in a more informative representation of the input conditions. For encoding the simulation time step $t_{as}$, we utilize Positional Encoding [65] followed by an MLP layer. This approach is specifically chosen to learn continuous representations that facilitate smooth transitions between consecutive time steps, thereby enhancing the model’s temporal coherence. In summary, this design offers the following advantages:
  • Algorithmic generalization. PICE improves the generalizability of the conditional diffusion model (see Table 2). The dual-encoding strategy leverages spatial information and captures non-linear relationships by incorporating both positional and polynomial encodings, empowering the model to adeptly manage a diverse array of simulation conditions, ranging from scenarios encountered during the training phase to conditions that lie beyond the spectrum of the training data.
  • Adaptable transferability. The PICE framework significantly augments our model’s adaptability to diverse physical scenarios and boundary conditions (see Table 2). By fine-tuning the pre-trained model with a minimal dataset, this methodology facilitates seamless adaptation, rendering our approach highly applicable across various domains within fusion research and other fields where precise and efficient simulations are imperative for deciphering intricate physical phenomena.
@@ 3.3 Rectified Flow-Based Acceleration
To enable rapid generation of high-fidelity synthetic data, we employ the Rectified Flow Acceleration (RFA) technique in model optimization. RFA leverages the principles of rectified flow [40, 18, 41] to convert the complex data trajectory from initial noise $\epsilon$ to the target snapshot $\mathcal{X}(t_{as})$ into a single streamlined denoising step, specifically aiming to distill the typically winding diffusion path into a direct and straight path — the shortest route between two distributions. During training, RFA minimizes the following objective function with straight ordinary differential equations:
$\operatorname*{arg\,min}_{\zeta}\quad\mathbb{E}\left[\int_{0}^{1}\left\|(\mathcal{X}(t_{as})-\epsilon)-\zeta(\mathcal{X}_{t},t~{}|~{}t_{as},\theta)\right\|^{2}dt\right],$ (7)
where $\mathcal{X}_{t}=t\mathcal{X}(t_{as})+(1-t)\epsilon$ denotes the linear interpolation between $\mathcal{X}(t_{as})$ and $\epsilon$ across the diffusion timeline, with $t$ ranging from 0 to 1. The score-based model $\zeta$, approximated using the U-Net, defines the learned trajectory. By minimizing the expectation of the squared deviations between the straight path $\mathcal{X}(t_{as})-\epsilon$ and the learned trajectory $\zeta(\mathcal{X}_{t},t~{}|~{}t_{as},\theta)$, RFA promotes the adoption of the shortest and most direct path in the transformation process, thus significantly reducing the denoising time. Once this time-dependent score-based model $\zeta$ is trained, we further straighten the learned trajectories through an interactive reflow procedure [41]. In summary, the RFA module provides additional benefits for our paradigm:
@@ 4.1 Experimental Setup
Datasets. We provide a novel dataset comprising 6,615 simulations across varied physical conditions, each containing 80 snapshots of electric fields along two orthogonal directions denoted E1 and E2. The data were generated by OSIRIS [19], a well-established PIC simulation software suite. The dataset covers diverse conditions, including Electron Temperature (Te), Ion Temperature (Ti), and Laser Intensity (Li), all of which are critical parameters influencing particle dynamics and the resultant electric fields. To foster further advancements in fusion research, we will release the dataset publicly upon acceptance.
Metrics. To evaluate the quality of the data synthesized by Diff-PIC, three metrics are adopted for comprehensive comparisons, including Fréchet Inception Distance (FID), Sliced Wasserstein Distance (SWD), and Maximal Mean Discrepancy (MMD). To validate physical soundness, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) are used to evaluate the energy difference between the synthetic physical environment and the ground truth.
Baselines. As there are few related studies to compare with, we evaluate two baselines as two design variations to demonstrate the effectiveness and speed of Diff-PIC. The first baseline (B-1), replaces the designed PICE with a simple MLP encoder, which is then integrated with the diffusion model. The second baseline (B-2), is a version of Diff-PIC that uses 1000 denoising steps through the DDPM noise scheduler [24]. These baselines help to highlight the contributions of the key components in the Diff-PIC architecture.
@@ 4.2 Main Results
In this section, we evaluate Diff-PIC on three key aspects for comprehensive comparisons. (1) The ability to interpolate and extrapolate data with novel physical conditions. (2) The physical validity of the synthetic physical environment. (3) The speedup and power efficiency compared to traditional PIC simulations. As a teaser, Fig. 3 compares the electric field results across 80 snapshots produced by Diff-PIC and PIC respectively in the testing set. The distributions on the right demonstrate that our proposed model successfully captures the data distributions in the ground truths.
@@
B-1 1.59 41.2 5.33e-4 1.87 46.8 7.21e-4 0.438 27.5 2.45e-4 0.513 37.6 2.89e-4
B-2 1.28 33.9 2.93e-4 1.54 44.3 5.11e-4 0.324 25.6 7.82e-5 0.336 35.2 9.13e-5
Diff-PIC 1.21 37.7 3.06e-4 1.62 40.4 5.35e-4 0.328 25.1 7.61e-5 0.341 34.3 8.98e-5
Interpolation and Extrapolation. The interpolation evaluation is carried out for a range of physical conditions (i.e., Te, Ti and Li). Among the conditions, 700 condition configurations are sampled, with 80% partitioned for training and the remaining 20% for testing. In Table 1, the remarkably low FID, SWD, and MMD scores indicate that the proposed Diff-PIC is able to synthesize high-quality scientific data similar to what PIC generates, meanwhile significantly outperforming baseline B-1 (i.e., without physically-informed conditional encoder equipped) on all three metrics. Compared to baseline B-2, the variations in scores are marginal, likely due to minor fluctuations in evaluation.
In terms of extrapolation, the capability of Diff-PIC is demonstrated with the electric field results on two orthogonal directions in Table 2 and Table 3. In the experiments, we progressively extend the range of physical conditions from the range selected for training. Specifically, the ranges are extended by 10%, 20%, 40%, and 80%. Similar to interpolation, the high quality remains for extrapolation, with considerable improvement from baseline B-1.
Table 2: Quantitative results for extrapolation evaluation on E1.
  Method 10% 20% 40% 80%
@@
B-1 3.92 81.7 3.27e-3 4.64 88.5 3.51e-3 5.25 91.0 4.36e-3 6.14 96.3 9.58e-3
B-2 3.51 78.3 2.64e-3 4.18 87.6 3.11e-3 4.74 89.4 3.96e-3 5.65 94.2 6.17e-3
Diff-PIC 3.87 81.2 2.76e-3 4.17 87.9 3.05e-3 4.96 90.5 3.91e-3 5.74 95.1 6.24e-3
Table 3: Quantitative results for extrapolation evaluation on E2.
  Method 10% 20% 40% 80%
[…]
B-1 1.94 47.3 7.72e-4 2.03 51.1 7.82e-4 2.19 60.8 8.64e-4 2.57 65.4 9.43e-4
B-2 1.73 44.3 7.28e-4 1.82 47.6 7.36e-4 2.01 52.2 7.83e-4 2.35 62.7 8.84e-4
Diff-PIC 1.81 47.5 7.34e-4 1.88 47.9 7.41e-4 1.93 51.9 7.76e-4 2.16 60.3 8.79e-4
Physical Validity. In addition to the high quality of the synthetic data, we test the physical validity of the generated electric fields at each snapshot with energy, a fundamental concept used to describe the general property of a physical system. In particular, the energy results of electric fields are illustrated in Fig. 4. We observe that the pattern of the ground truth is captured by Diff-PIC, with fairly low MAE and RMSE incurred. Especially, for electric field E2, the oscillation in energy is properly preserved in the synthetic data. Furthermore, this evaluation highlights that Diff-PIC is capable of generating sequentially continuous data, demonstrating the effectiveness of the proposed distillation paradigm and the physically-informed condition encoder.
@@
Speedup and Power Efficiency. In addition to the traditional PIC approach, B-1 and B-2 are also included in the comparison shown in Table 4. The PIC simulations are run on the Perlmutter supercomputer in the National Energy Research Scientific Computing (NERSC) facility, with AMD EPYC 7763 CPUs. As essential fusion phenomena typically appear at approximately 100 ps, the costs of PIC simulation at 100 ps are selected as the reference across the comparisons. The GPU results for Diff-PIC and baselines B-1 and B-2 are obtained on an Nvidia RTX 4090 GPU to demonstrate the availability of this approach to general users. The CPU results for these approaches are acquired on an Intel 13th Gen i9-13900KF CPU. The results highlight that Diff-PIC-GPU achieves over $10^{4}\times$ speedup versus traditional PIC simulation, as well as $10^{4}\times$ in terms of reduced energy cost.
@@
  Method PIC-100ps Diff-PIC-GPU Diff-PIC-CPU B-1-GPU B-1-CPU B-2-GPU B-2-CPU
Speedup 1.00$\times$ 1.62e4$\times$ 519$\times$ 1.66e4$\times$ 622$\times$ 9.63$\times$ 0.550$\times$
Energy Reduction 1.00$\times$ 1.01e4$\times$ 1.05e3$\times$ 1.03e4$\times$ 1.26e3$\times$ 6.04$\times$ 1.11$\times$
@@ 4.3 Discussion
To provide further insights and highlight the value of this work, it is worth noting that our approach exhibits outstanding scalability compared to traditional PIC simulations. In PIC, for $N$ particles, the computing complexity can reach a formidable $N^{2}$, as each particle interacts with every other particle. On the contrary, Diff-PIC focuses on the macroscopic data (e.g., electric field) in the form of distribution, therefore it is not sensitive to the number of particles in space. For larger particle spaces, the speedup achieved by Diff-PIC can be easily improved by extra orders of magnitude, further accelerating the research of fusion, or other research areas involving large-scale PIC simulations.
@@ 5 Related Work
Contrastingly, our proposed method overcomes these limitations by employing a conditional diffusion model to distill the complex patterns captured by PIC simulations from a limited training dataset. Utilizing a time-dependent score-based model, our approach can efficiently generate high-fidelity synthetic data (see Fig. 3 and Table 1) without the computational expense of traditional PIC algorithms (see Table 4). This results in a significant reduction in computational cost while maintaining high simulation accuracy. Moreover, our method is highly adaptable, as it can be fine-tuned with minimal additional data to suit various physical conditions (see Table 2 and 3). This flexibility renders our approach suitable for a wide range of applications in fusion research and beyond, where efficient and accurate simulations are crucial for understanding complex physical phenomena.

### v3

@@ 3.2 Physically-Informed Parameter Encoder
PIC simulations employ continuous physical parameters as inputs, which necessitate a seamless and continuous transition in the resulting synthesized data as input parameters are adjusted. Consequently, an encoder is considered exceptionally beneficial in this scenario, responsible for transforming domain-specific inputs into embeddings comprehensible by the model, meanwhile preserving the physical continuity of output snapshots. In particular, these inputs comprise the simulation parameters $\theta$ and the target simulation snapshot $t_{as}$ that the conditional diffusion model aims to generate.
@@
Given the extensive range of physically feasible parameters and the limited data available during training, an optimal encoder must excel in both interpolation and extrapolation — critical measures of the model’s ability of generalization. Interpolation capability refers to the encoder’s proficiency in generating suitable embeddings for new parameters that, although not encountered during training, lie between observed parameters. Extrapolation capability, conversely, pertains to generating embeddings for parameters that fall outside the range of those observed during training. Notably, both capabilities are indispensable for addressing the LPI problem, in order to cover a large enough, and fine-grained enough parameter space to provide sufficient insight into further LPI evaluations.
Figure 3: The proposed PIPE.
@@
To meet the two essentials, we introduce a Physically-Informed Parameter Encoder (PIPE) as shown in Fig. 3. To encode the simulation parameters $\theta$, we incorporate two distinct types of encoders. For interpolation, we employ Positional Encoding [57] (denoted $\sim$” in a circle) to learn a continuous representation of input parameters, enabling smooth transitions between observed parameters. To augment extrapolation capability, we enhance the encoding process with a polynomial encoder (“P” in a circle), utilizing transformation functions constructed as a linear combination of polynomial basis functions $f_{i}(\theta)$ of varying degrees:
@@
where $n$ denotes the maximum order of polynomial terms, and the polynomial $\mathcal{P}(\theta)$ can be chosen as Chebyshev polynomials and Legendre polynomials, based on the characteristics of the parameter space. This polynomial enhancement allows the encoder to generate plausible embeddings for parameters well beyond those encountered during training, ensuring robust performance across a broader spectrum of simulation scenarios. Subsequently, we concatenate (“C” in a circle) the embeddings from these two encoders and apply a Multi-Layer Perceptron (MLP) to further refine the embeddings. The MLP, with its trainable parameters, learns to combine and transform the concatenated embeddings, resulting in a more informative representation of the input parameters. For encoding the simulation time step $t_{as}$, we utilize Positional Encoding  [57] followed by an MLP layer. This approach is specifically chosen to learn continuous representations that facilitate smooth transitions between consecutive snapshots, thereby enhancing the model’s temporal coherence. In summary, this design offers the following advantages:
  • Algorithmic generalization. PIPE improves the generalizability of the conditional diffusion model (see Table 2). The dual-encoding strategy captures non-linear relationships by incorporating both positional and polynomial encodings, empowering the model to adeptly manage a diverse array of simulation parameters, ranging from scenarios encountered during the training phase to parameters that lie beyond the spectrum of the training data.
  • Adaptive transferability. By fine-tuning the pre-trained model with a new dataset, this methodology facilitates adaptation to other fields where precise and efficient simulations are imperative for deciphering intricate physical phenomena.
@@ 3.3 Rectified Flow-Based Acceleration
To enable rapid generation of high-fidelity synthetic data, we employ the Rectified Flow Acceleration (RFA) technique in model optimization. Based on the principles of rectified flow [38, 16, 39], RFA converts the original complex denoising trajectory from initial noise $\epsilon$ to the target electric field snapshot $\mathbf{E}(t_{as})$ into a direct and straight path – the shortest route between two distributions. During training, RFA minimizes the following objective function with straight ordinary differential equations:
$\operatorname*{arg\,min}_{\zeta}\quad\mathbb{E}\left[\int_{0}^{1}\left\|(\mathbf{E}(t_{as},\theta)-\epsilon)-\zeta(\mathbf{E}_{t},t~{}|~{}t_{as},\theta)\right\|^{2}dt\right],$ (5)
where $\mathbf{E}_{t}=t\mathbf{E}(t_{as},\theta)+(1-t)\epsilon$ denotes the linear interpolation between $\mathbf{E}(t_{as},\theta)$ and $\epsilon$ across the diffusion timeline, with $t$ ranging from 0 to 1. The score-based model $\zeta$, approximated using the U-Net, defines the learned trajectory. By minimizing the expectation of the squared deviations between the straight path $\mathbf{E}(t_{as},\theta)-\epsilon$ and the learned trajectory $\zeta(\mathbf{E}_{t},t~{}|~{}t_{as},\theta)$, RFA promotes the adoption of the shortest and most direct path in the denoising process, thus significantly reducing the denoising time. Once this time-dependent score-based model $\zeta$ is trained, we further straighten the learned trajectories through an interactive reflow procedure  [39].
In summary, the RFA module provides additional benefits for our paradigm:
@@ 4.1 Experimental Setup
Datasets. We provide a new dataset comprising 6,615 simulations across varied physical simulation parameters, each containing 80 snapshots of electric fields along two orthogonal directions denoted E1 and E2. The data were generated by OSIRIS [17], a well-established PIC simulation software suite. The dataset covers diverse parameters, including $T_{e}$, $T_{i}$, and $I$, all of which are critical parameters influencing the resultant electric fields. To foster further advancements in fusion and scientific data generation research, we will release the dataset publicly.
Metrics. To validate physical soundness, Mean Absolute Error (MAE) and Root Mean Squared Error (RMSE) are used to evaluate the electric field difference and the energy difference between the Diff-PIC generated and the ground truth produced by PIC simulations. To better demonstrate the relative error, the dataset is normalized to [0,1]. To further evaluate the difference in the generated and the ground truth data distributions, the Fréchet Inception Distance (FID) metric is also employed, reflecting the fidelity of the electric fields produced by DiffPIC.
Baselines. We compare Diff-PIC with two other SOTA generative models, Generative Adversarial Networks [26] and Normalizing Flow [62]. The baseline models are implemented based on the setups provided in the original papers. Since neither of them originally supports learning meaningful embeddings for the physical parameters, for fair comparison, we equip them with the proposed PIPE to establish two baselines: GAN-PIC and NF-PIC.
Diff-PIC configurations. The architectural foundation of our model is predicated on the U-Net framework [43], incorporating a series of three down-sampling blocks followed by three up-sampling blocks. The training regimen of our model encompasses an extensive 600 epochs, employing a batch size of 64, a configuration empirically validated to secure model convergence. Furthermore, the training protocol adheres to a fixed learning rate of $5\times 10^{-4}$, optimized via the Adam optimizer [28]. The optimization objective is delineated in Eq. 5. In addition, in terms of the encoders, the positional encoding mechanism yields a 16-dimensional embedding from a single input, while the polynomial encoder operates with a highest polynomial order of 4.
@@ 4.2 Main Results
In this section, we evaluate Diff-PIC on three key aspects for comprehensive comparisons. (1) The interpolation ability and extrapolation ability. (2) The physical validity of the Diff-PIC generated data. (3) The speedup and power efficiency compared to traditional PIC simulations.
@@
GAN-PIC 4.59e-2 5.31e-2 2.32 4.73e-2 5.84e-2 2.51 1.82e-2 2.07e-2 0.973 1.97e-2 2.18e-2 1.03
NF-PIC 4.46e-2 5.12e-2 2.06 4.61e-2 5.35e-2 2.42 1.70e-2 2.03e-2 0.914 1.86e-2 2.45e-2 0.986
Diff-PIC 1.56e-2 2.67e-2 1.21 1.68e-2 2.29e-2 1.62 7.95e-3 9.32e-3 0.328 8.26e-3 1.03e-2 0.341
Interpolation and Extrapolation. To evaluate the interpolation capability of Diff-PIC, we sample a specified range for each simulation parameters ($T_{e}$, $T_{i}$, and $I$), totally 500 simulations and $500\times 80=40,000$ snapshots. Then, we randomly split these 500 simulations into training and testing set with the ratio of 80% and 20%. We train Diff-PIC and baselines on the training set, and report the performance of Diff-PIC on the training and testing set in Table 1. The reasonably low MAE, RMSE, and FID scores indicate that the proposed Diff-PIC is able to synthesize high-quality scientific data similar to what PIC generates, meanwhile significantly outperforming baselines on all three metrics. On average, Diff-PIC achieves 59.25% reduced MAE, 57.77% reduced RMSE, and 49.21% reduced FID for both testing sets, compared to the baselines.
In addition, Fig. 4 compares the results for one randomly selected simulation produced by Diff-PIC and PIC respectively in the testing set. Throughout the snapshots ($t_{as}=1\rightarrow 30\rightarrow 50\rightarrow 80$ for example), the distributions of the synthetic data closely follow the ground truths, indicating that physical continuity is maintained over time. Additionaly, the distributions on the right demonstrate that our proposed model successfully captures the data distributions in the ground truths.
In terms of extrapolation capability evaluation, we progressively extend the range of physical parameters from the range selected for training. In particular, the ranges are extended by 10% and 20% respectively as case studies. The results on Table 2 demonstrate that Diff-PIC achieves approximately only 2% relative absolute error in extrapolation tasks, significantly outperforming the other generative counterparts. On average, Diff-PIC achieves 59.16% reduced MAE, 56.53% reduced RMSE, and 29.70% reduced FID for all test cases, compared to the baselines.
Table 2: Quantitative results for extrapolation evaluation.
  Method E1 10% E1 20% E2 10% E2 20%
@@
GAN-PIC 5.24e-2 6.32e-2 1.97 5.73e-2 6.98e-2 2.15 2.18e-2 3.25e-2 1.15 2.93e-2 3.89e-2 1.29
NF-PIC 4.74e-2 5.31e-2 1.85 5.41e-2 6.46e-2 2.08 1.90e-2 3.16e-2 1.04 2.63e-2 3.42e-2 1.17
Diff-PIC 1.83e-2 2.40e-2 1.74 2.18e-2 2.62e-2 1.82 9.47e-3 1.36e-2 0.536 1.13e-2 1.85e-2 0.673
Physical Validity. In addition to the high quality of the synthetic data, we test the physical validity of the generated electric fields at each snapshot with energy, a fundamental concept used to describe the general property of a physical system. In particular, we randomly pick a simulation in testing set, the energy results of its electric fields are illustrated in Fig. 5. We observe that the pattern of the ground truth is captured by Diff-PIC, with fairly low MAE and RMSE incurred. Especially, for electric field E2, the oscillation in energy is properly preserved in the synthetic data. Furthermore, this evaluation highlights that Diff-PIC is capable of generating sequentially continuous data, demonstrating the effectiveness of the proposed distillation paradigm and the Physically-Informed Parameter Encoder.
@@
  Method PIC-100 ps Diff-PIC-GPU Diff-PIC-CPU GAN-PIC-GPU GAN-PIC-CPU NF-PIC-GPU NF-PIC-CPU
Speedup 1.00$\times$ 1.62e4$\times$ 519$\times$ 1.56e4$\times$ 523$\times$ 9.21e2$\times$ 24$\times$
Energy Reduction 1.00$\times$ 1.01e4$\times$ 1.05e3$\times$ 1.13e4$\times$ 1.47e3$\times$ 8.14e2$\times$ 53$\times$
Speedup and Power Efficiency. In addition to the traditional PIC approach, baselines are also included in the comparison shown in Table 3. The PIC simulations are run on the Perlmutter supercomputer in the National Energy Research Scientific Computing (NERSC) facility, with AMD EPYC 7763 CPUs. As essential fusion phenomena typically appear at approximately 100 ps, the costs of PIC simulation at 100 ps are selected as the reference across the comparisons. The GPU results for Diff-PIC and baselines are obtained on an Nvidia RTX 4090 GPU to demonstrate the availability of this approach to general users. The CPU results for these approaches are acquired on an Intel 13th Gen i9-13900KF CPU. The results highlight that Diff-PIC-GPU achieves over $10^{4}\times$ speedup versus traditional PIC simulation, as well as $10^{4}\times$ in terms of reduced energy cost.
@@ 4.3 Discussion
To provide further insights and highlight the value of this work, it is worth noting that our approach exhibits outstanding scalability compared to traditional PIC simulations. In PIC, for $N$ particles, the computing complexity can generally reach a formidable $TN\text{log}N$ with $T$ time steps. In contrast, Diff-PIC is not sensitive to the number of particles in space nor the number of time steps, since it focuses on generating the macroscopic data (e.g., electric field, which is usually considered more informative than individual particle status) directly for a specific time. For larger particle spaces, the speedup achieved by Diff-PIC can be easily improved by extra orders of magnitude, further accelerating the research of fusion, or other research areas involving large-scale PIC simulations.
@@ 5 Related Work
Contrastingly, our proposed method overcomes these limitations by employing a conditional diffusion model to distill the complex patterns captured by PIC simulations from a limited training dataset. Utilizing a time-dependent score-based model, our approach can efficiently generate high-fidelity synthetic data (see Fig. 4 and Table 1) without the computational expense of traditional PIC algorithms (see Table 3). This results in a significant reduction in computational cost while maintaining high simulation accuracy. Moreover, our method is highly adaptable, as it can be fine-tuned with minimal additional data to suit various physical parameters. This flexibility renders our approach suitable for a wide range of applications in fusion research and beyond, where efficient and accurate simulations are crucial for understanding complex physical phenomena.